All of the carbon dioxide in a can of lemonade with an initial mass of 381.04g was carefully extracted and collected as a gas. The final mass of the drink was 380.41g. What volume would the carbon dioxide occupy at 100kPa and 25oC?
1
Expert's answer
2020-10-12T13:29:25-0400

m(CO2) = minitial - mfin al = 381.04 - 380.41 = 0.63 g

p = 100 kPa = 100000 Pa

T = t+273 = 25+273 = 298

M = 44 g/mol

pV = (m/M)*RT

V = "\frac{mRT}{Mp}"

R = 8.314 J/(mol*K)

"V = \frac{0.63*8.314*298}{44*100000} = 3.55 * 10^{-4}\ (m^{3})"